what is the difference between -2 to the power of 2
and (-2) to the power of 2
: what is the difference between -2 to the power of 2
and (-2) to the power of 2

Answer by checkley77(3848)

-(2)^2=-(2*2)=-4
(-2)^2=(-2*-2)=4